### v1

# What is a Schiff moment anyway?

Abstract
Schiff moments of atomic nuclei are of considerable interest to experiments searching for undiscovered new physics that breaks time-reversal symmetry. I develop a simple picture of the Schiff moment of a charge distribution, and discuss the interaction of the Schiff moment of a nucleus with the field produced by an electron in an atom.
Experimental searches for the violation of time-reversal (T) symmetry in nuclei typically focus on a property known as the nuclear Schiff moment [1, 2]. Currently, a host of experiments are searching for Schiff moments of various nuclei using precision techniques, in systems as varied as atoms [3, 4], molecules [5] and solids [6], hoping to find clues about new sources of T-violation that could explain the origin of the matter-antimatter asymmetry of the universe [7]. So this small nuclear moment has lofty implications.
But what is a Schiff moment? Informally, a Schiff moment is sometimes described as the difference between the mean-squared radii of the charge and the electric-dipole distributions [8, 9]. Formal definitions can be found in a number of technical articles, e.g., [10, 11, 12, 13, 14] – however, these are often written in the sophisticated language of theoretical nuclear physics, which can obscure the simple electrostatics underlying the notion of a Schiff moment.
The purpose of this note is two-fold: first, to provide a simple yet accurate description of the Schiff moment of a charge distribution, and second, to illustrate the interesting physics of the interaction of a nuclear Schiff moment with electrons in atoms. This note may be useful to undergraduates or graduate students interested in precision searches for physics beyond the Standard Model of particle physics. Perhaps it may be useful as well to students and instructors of electromagnetism courses, as an interesting example of basic electrostatics that can be connected to contemporary research.
I.   Electron interaction with a nuclear charge distribution
The potential energy for the electrostatic interaction between an electron (position $\vec{r}$, charge e-e) and a nuclear charge distribution (position $\vec{R}$, charge density $\rho_{n}$) is
$V_{en}(\vec{r})=\frac{-e}{4\pi\epsilon_{0}}\int d^{3}R\,\frac{\rho_{n}(\vec{R})}{|\vec{r}-\vec{R}|}.$ (1)
The integral $\int d^{3}R$ is over the volume of the nucleus. To break up this interaction into simpler parts, consider the Taylor expansion of $1/|\vec{r}-\vec{R}|$ that appears inside the integral.
$\frac{1}{|\vec{r}-\vec{R}|}=\frac{1}{r}-X_{i}\,\nabla_{i}\frac{1}{r}+\frac{1}{2!}X_{i}X_{j}\,\nabla_{j}\nabla_{i}\frac{1}{r}-\frac{1}{3!}X_{i}X_{j}X_{\ell}\,\nabla_{\ell}\nabla_{j}\nabla_{i}\frac{1}{r}+\ldots$ (2)
Here $X_{i}$ are the components of $\vec{R}$, $x_{i}$ are the components of $\vec{r}$ and $\nabla_{i}=\frac{\partial}{\partial x_{i}}$. The cartesian indices $i,j,\ell$ range from 1 to 3. Repeated indices are summed over.
The Taylor expansion allows us to separate the electron coordinate and its derivatives from the integral over $d^{3}R$. As a result, the electron-nucleus interaction conveniently splits into a sum of terms, each of which is a product of a nuclear quantity and an electronic quantity. Up to 3rd order in $\vec{R}$, the interaction potential is $V_{en}=V_{E0}+V_{E1}+V_{E2}+V_{E3}$, where
$\begin{split}V_{E0}&=k\left(\int d^{3}R\,\rho_{n}\right)\left[\frac{1}{r}\right(ht]\\ V_{E1}&=-k\left(\int d^{3}R\,\rho_{n}\,X_{i}\right)\left[\frac{-x_{i}}{r^{3}}\right]\\ V_{E2}&=\frac{k}{2}\left(\int d^{3}R\,\rho_{n}\,X_{i}X_{j}\right)\left[-\frac{4\pi}{3}\delta_{ij}\,\delta(\vec{r})+\frac{3x_{i}x_{j}-r^{2}\delta_{ij}}{r^{5}}\right]\\ V_{E3}&=-\frac{k}{6}\left(\int d^{3}R\,\rho_{n}\,X_{i}X_{j}X_{\ell}\right)\left[-\frac{4\pi}{5}\delta_{\{ij}\,\nabla_{\ell\}}\delta(\vec{r})+\frac{3\delta_{\{ij}x_{\ell\}}r^{2}-15x_{i}x_{j}x_{\ell}}{r^{7}}\right].\end{split}$ (3)
Here $k=-e/4\pi\epsilon_{0}$ for ease of notation. The $E0,E1,E2$ and $E3$ interactions are traditionally called the electric monopole, dipole, quadrupole and octupole interactions respectively. Each of these is a product of a nuclear multipole moment (in round brackets) that depends on $\vec{R}$, and a field produced by the electron (in square brackets) that depends on $\vec{r}$. The gradients of $1/r$ appearing above are derived in the Appendix. The symmetrized 3rd rank tensor formed from the 2nd rank tensor $A_{ij}$ and the vector $B_{\ell}$ is compactly written as $A_{\{ij}B_{\ell\}}=A_{ij}B_{\ell}+A_{\ell i}B_{j}+A_{j\ell}B_{i}$.
In the $E2$ and $E3$ interactions, there are two kinds of fields produced by the electron. The so-called “contact” terms contain $\delta(\vec{r})$ or its derivatives [15, 16], whereas the non-contact terms are traceless tensors. Let us focus on the contact $E3$ term
$\begin{split}V_{cE3}&=-\frac{e}{30\epsilon_{0}}\left(\int d^{3}R\,\rho_{n}(\vec{R})\,X_{i}X_{j}X_{\ell}\right)\left[(\delta_{ij}\nabla_{\ell}+\delta_{j\ell}\nabla_{i}+\delta_{\ell i}\nabla_{j})\delta(\vec{r})\right]\\ &=-\left(\frac{1}{10}\int d^{3}R\,\rho_{n}(\vec{R})\,R^{2}\,\vec{R}\right)\cdot\left[\frac{e}{\epsilon_{0}}\vec{\nabla}\delta(\vec{r})\right]\\ &=-\vec{\mathscr{S}}_{0}\cdot\vec{\mathcal{F}}.\end{split}$ (4)
The quantity $\vec{\mathscr{S}}_{0}=\frac{1}{10}\int d^{3}R\,\rho_{n}(\vec{R})\,R^{2}\,\vec{R}$ is the first piece of the Schiff moment of the charge distribution $\rho_{n}(\vec{R})$. I have also defined $\vec{\mathcal{F}}=\frac{e}{\epsilon_{0}}\vec{\nabla}\delta(\vec{r})$, the field from the electron that interacts with the nuclear Schiff moment – this will be discussed in more detail in Section IV.
But there is more to the story: we need to consider one more effect for a nucleus in an atom.
II.   Shielding of the nuclear electric dipole moment in an atom
An electric dipole moment of a charge distribution that also has nonzero total charge is equivalent to a displacement of the centre of charge. Inside an atom, this fact has an interesting consequence.
Let the nucleus have a nonzero electric dipole moment (EDM) $\vec{D}=\int d^{3}R\,\rho_{n}\vec{R}$. The combination of its electric charge $Q=\int d^{3}R\rho_{n}$ (monopole moment) and dipole moment implies that the centre of charge of the nucleus is at $\Delta\vec{R}=\frac{\vec{D}}{Q}$.
In atomic physics, electron wavefunctions are usually calculated with respect to an origin located at the centre of charge of the nucleus. Therefore, physically displacing the nuclear centre of charge by a small amount $\Delta\vec{R}$ – sufficiently slowly for the adiabatic theorem [17] to apply – just displaces all the electron eigenfunctions by $\Delta\vec{R}$ as well. This operation is mathematically the same as shifting the origin to the centre of charge of the nucleus, which is equivalent to replacing $\vec{r}$ with $\vec{r}+\Delta\vec{R}$ in Equation (3).
The modified electron-nuclear interaction (which I will denote as $V_{en}^{\prime}$), up to first order in $\vec{D}$, is
$V^{\prime}_{en}(\vec{r})=V_{en}(\vec{r}+\Delta\vec{R})=V_{en}(\vec{r})+\frac{D_{i}}{Q}\nabla_{i}V_{en}(\vec{r})=V_{E0}^{\prime}+V_{E1}^{\prime}+V_{E2}^{\prime}+V_{E3}^{\prime},$ (5)
where the modified multipole interactions are
$\begin{split}V^{\prime}_{E0}&=k\left(\int d^{3}R\,\rho_{n}\right)\left[\frac{1}{r}\right]\\ V^{\prime}_{E1}&=-k\left(\int d^{3}R\,\rho_{n}X_{i}\right)\left[-\frac{x_{i}}{r^{3}}\right]+\frac{D_{j}}{Q}\nabla_{j}V_{E0}\\ V^{\prime}_{E2}&=\frac{k}{2}\left(\int d^{3}R\,\rho_{n}\,X_{i}X_{j}\right)\left[-\frac{4\pi}{3}\delta_{ij}\,\delta(\vec{r})+\frac{3x_{i}x_{j}-r^{2}\delta_{ij}}{r^{5}}\right]+\frac{D_{i}}{Q}\nabla_{i}V_{E1}\\ V^{\prime}_{E3}&=-\frac{k}{6}\left(\int d^{3}R\,\rho_{n}\,X_{i}X_{j}X_{\ell}\right)\left[-\frac{4\pi}{5}\delta_{\{ij}\nabla_{\ell\}}\delta(\vec{r})+\frac{3\delta_{\{ij}x_{\ell\}}r^{2}-15x_{i}x_{j}x_{\ell}}{r^{7}}\right]+\frac{D_{i}}{Q}\nabla_{i}V_{E2}.\end{split}$ (6)
First, consider the $E1$ interaction, which becomes
$\begin{split}V^{\prime}_{E1}&=-k\left(\int d^{3}R\,\rho_{n}X_{i}\right)\left[-\frac{x_{i}}{r^{3}}\right]+k\frac{D_{j}}{Q}\left(\int d^{3}R\,\rho_{n}\right)\left[-\frac{x_{j}}{r^{3}}\right]=0,\end{split}$ (7)
since $Q=\int d^{3}R\,\rho_{n}$ and $D_{i}=\int d^{3}R\,\rho_{n}X_{i}$. The EDM of the nucleus is shielded by the rearrangement of the electron charge distribution, leading to zero measurable contribution of the EDM to the $E1$ interaction. This is Schiff’s shielding theorem [1].
Next consider the contact part of the modified $E3$ interaction, which acquires a correction from the original contact $E2$ interaction to become
$\begin{split}V^{\prime}_{cE3}&=-\left(\frac{1}{10}\int d^{3}R\,\rho_{n}R^{2}X_{i}\right)\cdot\left[\frac{e}{\epsilon_{0}}\nabla_{i}\delta(\vec{r})\right]+\frac{e}{6\epsilon_{0}}\frac{D_{j}}{Q}\left(\int d^{3}R\,\rho_{n}R^{2}\right)\left[\nabla_{j}\delta(\vec{r})\right]\\ &=-\left(\frac{1}{10}\int d^{3}R\,\rho_{n}R^{2}X_{i}-\frac{1}{6}\frac{D_{i}}{Q}\int d^{3}R\,\rho_{n}R^{2}\right)\left[\frac{e}{\epsilon_{0}}\nabla_{i}\delta(\vec{r})\right]\\ &=-(\vec{\mathscr{S}}_{0}+\vec{\mathscr{S}}_{1})\cdot\vec{\mathcal{F}}.\end{split}$ (8)
The combination
$\vec{\mathscr{S}}=\vec{\mathscr{S}}_{0}+\vec{\mathscr{S}}_{1}=\frac{1}{10}\int d^{3}R\,\rho_{n}(\vec{R})R^{2}\vec{R}-\frac{1}{6}\frac{\vec{D}}{Q}\int d^{3}R\,\rho_{n}(\vec{R})R^{2}$ (9)
is the Schiff moment of the nucleus. It consists of the nuclear moment $\vec{\mathscr{S}}_{0}$, defined in Equation (4), plus a correction, $\vec{\mathscr{S}}_{1}\propto\vec{D}$, due to the displacement of the centre of charge.
The Schiff moment $\vec{\mathscr{S}}$ is purely a property of the distribution $\rho_{n}(\vec{R})$. Some of its properties are:
  1. (i)
    $\vec{\mathscr{S}}$ can be defined for any charge distribution, even in classical electromagnetism. There is nothing intrinsically nuclear or quantum-mechanical about it.
  2. (ii)
    $\vec{\mathscr{S}}$ has units of charge $\times$ length3 (C m3 in SI, or $e$ fm3 in units appropriate for nuclei).
  3. (iii)
    The Schiff moment is a vector quantity. But it is not the EDM: the contact $E3$ interaction is not the same as the $E1$ interaction.
  4. (iv)
    In a quantum mechanical system with angular momentum, such as a nucleus with spin $\vec{I}$, the Schiff moment is parallel to the total angular momentum, $\vec{\mathscr{S}}\propto\vec{I}$. This property follows from the Wigner-Eckart theorem [18].
III.   Charge distribution with a Schiff moment
For concreteness, here is a simple example of a charge distribution with a nonzero Schiff moment. Consider a charge distribution confined to the surface of a sphere of radius $b$. Let the charge per unit area on the surface be $\sigma(\theta,\phi)=\frac{e}{4\pi b^{2}}\left(1+\eta\cos\theta\right)$, where $\theta,\phi$ are the usual spherical polar angles. Let us assume the dimensionless constant $\eta\ll 1$. The moments of this distribution are
$\begin{split}\int d^{3}R\,\rho(\vec{R})&\to b^{2}\int_{0}^{2\pi}d\phi\int_{0}^{\pi}\sin\theta\,d\theta\,\sigma(\theta,\phi)=e\\ \int d^{3}R\,\rho(\vec{R})\,Z&\to b^{3}\int_{0}^{2\pi}d\phi\int_{0}^{\pi}\sin\theta\,d\theta\,\sigma(\theta,\phi)\,\cos\theta=\frac{1}{3}\eta eb\\ \int d^{3}R\,\rho(\vec{R})\,R^{2}&\to b^{4}\int_{0}^{2\pi}d\phi\int_{0}^{\pi}\sin\theta\,d\theta\,\sigma(\theta,\phi)=eb^{2}\\ \int d^{3}R\,\rho(\vec{R})\,R^{2}Z&\to b^{5}\int_{0}^{2\pi}d\phi\int_{0}^{\pi}\sin\theta\,d\theta\,\sigma(\theta,\phi)\,\cos\theta=\frac{1}{3}\eta eb^{3}.\end{split}$ (10)
Since the charge distribution has azimuthal symmetry, the $X$ and $Y$ components of the EDM and Schiff moment are zero, and only the $Z$ components are nonzero. This distribution has both a monopole and a dipole moment, and so the centre of charge is not at the centre of the sphere, but at $\Delta\vec{R}=(0,0,\frac{\eta b}{3})$ instead. The $Z$ component of $\vec{\mathscr{S}}$ is easily calculated using Eq. (9) to be $\mathscr{S}_{Z}=\frac{1}{30}\eta eb^{3}-\frac{1}{18}\eta eb^{3}=-\frac{1}{45}\eta eb^{3}$.
IV.   The $\mathcal{F}$-field
Having seen how the definition of the Schiff moment arises naturally from the electron-nucleus electrostatic interaction, let us now consider the electron-generated field that couples to the nuclear Schiff moment. The definition of $\vec{\mathcal{F}}$ in Equation (8), which involves the gradient of a delta-function, is a rather singular mathematical object. To gain physical understanding without getting lost in mathematical subtleties, let us replace the $\delta(\vec{r})$ in the definition of $\vec{\mathcal{F}}$ with its smooth regularized version described in the Appendix, with a regularization length scale $a$. The gradient of the delta-function is then
$\begin{split}\vec{\nabla}\delta(\vec{r})&\approx\frac{3}{4\pi a^{3}}\vec{\nabla}\left(1+\frac{r^{2}}{a^{2}}\right)^{-5/2}\\ &=-\frac{15}{4\pi a^{5}}\frac{\vec{r}}{\left(1+\frac{r^{2}}{a^{2}}\right)^{7/2}}\\ &=-\frac{\vec{r}\,g(r)}{a^{5}},\end{split}$ (11)
where the dimensionless function $g(r)=\frac{15}{4\pi}\left(1+\frac{r^{2}}{a^{2}}\right)^{-7/2}$ is defined for convenience. Thus a finite approximation of the $\mathcal{F}$-field produced by the electron is $\vec{\mathcal{F}}=-\frac{e\vec{r}\,g(r)}{\epsilon_{0}a^{5}}$. Fig. 1 shows an illustration of the $\mathcal{F}$-field from an electron in a classical orbit around a nucleus.
Figure 1: The $\mathcal{F}$-field produced by an electron in an atom. This field is localized within a sphere of radius $\sim a$ (which we take to be much smaller than the size of an electron orbit), and is directed opposite to the electron’s position vector $\vec{r}$. The $\mathcal{F}$-field couples to the Schiff moment of the nucleus as described by Equation (8).
But electrons do not move classically inside atoms, so it is instructive to calculate the $\mathcal{F}$-field produced by an electron in a hydrogenic atom. Consider an electron in a superposition of nsns and $n^{\prime}p_{z}$ states, where $n,n^{\prime}$ denote hydrogenic principal quantum numbers: ${\left|{\psi}\right\rangle}=c_{s}{\left|{ns}\right\rangle}+c_{p}{\left|{n^{\prime}p_{z}}\right\rangle}$, with complex coefficients $c_{s}$ and $c_{p}$. The spatial wavefunctions of the atomic states are ${\left\langle{\vec{r}}|{ns}\right\rangle}=\mathfrak{R}_{ns}(r)Y_{00}(\theta\phi)$ and ${\left\langle{\vec{r}}|{n^{\prime}p_{z}}\right\rangle}=\mathfrak{R}_{n^{\prime}p}(r)Y_{10}(\theta\phi)$, where $\mathfrak{R}_{n\ell}(r)$ are hydrogenic radial functions and $Y_{\ell m}(\theta\phi)$ are spherical harmonics. Note that ${\left\langle{ns}\right|}\mathcal{F}_{z}{\left|{ns}\right\rangle}$ and ${\left\langle{n^{\prime}p}\right|}\mathcal{F}_{z}{\left|{n^{\prime}p}\right\rangle}$ both vanish, since $\mathcal{F}_{z}$ is an odd-parity operator sandwiched between definite-parity wavefunctions, which makes the expectation value integral equal zero. Therefore only the cross terms between $\left|{ns}\right\rangle$} and $\left|{n^{\prime}p}\right\rangle$} contribute, and so ${\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}=2\,\mathrm{Re}(c_{p}^{*}c_{s}){\left\langle{n^{\prime}p}\right|}\mathcal{F}_{z}{\left|{ns}\right\rangle}$. Defining $\xi=\mathrm{Re}(c_{p}^{*}c_{s})$, the expectation value of the $z$-component of $\vec{\mathcal{F}}$ in the electron state $\left|{\psi}\right\rangle$} is
$\begin{split}{\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}&=2\xi{\left\langle{n^{\prime}p}\right|}\mathcal{F}_{z}{\left|{ns}\right\rangle}\\ &=-\frac{2\xi e}{\epsilon_{0}a^{5}}\int_{0}^{\infty}4\pi r^{2}dr\,\mathfrak{R}_{ns}(r)\mathfrak{R}_{n^{\prime}p}(r)\,rg(r)\int_{0}^{2\pi}d\phi\int_{0}^{\pi}d\theta\,\sin\theta\,Y_{00}(\theta\phi)\cos\theta\,Y_{10}(\theta\phi)\\ &=-\frac{2\xi e}{\sqrt{3}\epsilon_{0}a^{5}}\int_{0}^{r_{1}}4\pi r^{2}dr\,\mathfrak{R}_{ns}(r)\mathfrak{R}_{n^{\prime}p}(r)\,rg(r).\end{split}$ (12)
In the last line, I also used the fact that $g(r)$ drops rapidly to zero when $r$ is much larger than $a$. So it is sufficient to take the radial integral to an upper limit $r_{1}\gtrsim a$ (say $r_{1}=5a$), instead of $\infty$. The exact value of $r_{1}$ will turn out to be unimportant.
Now imagine that we set the length scale $a$ to be much smaller than the typical size of an atom, so that inside a sphere of radius $a$ we can approximate $\mathfrak{R}_{ns}(r)\approx\alpha$ and $\mathfrak{R}_{n^{\prime}p}(r)=\beta r$, their values close to the origin. The constants $\alpha$ and $\beta$ can be obtained from the leading-order series approximation of hydrogenic wavefunctions, and their values are $\alpha=2\sqrt{\frac{Z^{3}}{n^{3}a_{0}^{3}}}$ and $\beta=\frac{2Z}{3a_{0}}\sqrt{\frac{Z^{3}}{n^{\prime 3}a_{0}^{3}}\left(1-\frac{1}{n^{\prime 2}}\right)}$, where $Z$ is the nuclear charge and $a_{0}$ is the Bohr radius.
Thus the $\mathcal{F}$-field produced by the electron is
$\begin{split}{\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}=-\frac{2\xi e}{\sqrt{3}\epsilon_{0}a^{5}}\alpha\beta\int_{0}^{r_{1}}4\pi r^{2}dr\,r^{2}g(r).\end{split}$ (13)
Remarkably, the integral $\int_{0}^{r_{1}}4\pi r^{2}dr\,r^{2}g(r)=3a^{5}$ whenever $r_{1}\gtrsim a$, since the step-like function $g(r)$ only has appreciable support over a region of radius $\sim a$ around the origin. Hence ${\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}$ has a finite and well-defined value, regardless of our choice of specific values for the regularization length scale $a$ and the integral limit $r_{1}$ as long as $a\ll r_{1}\ll a_{0}/Z$. Thus
${\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}=-\frac{2\sqrt{3}\xi e}{\epsilon_{0}}\alpha\beta=-\frac{8\xi}{\sqrt{3}}\frac{eZ^{4}}{\epsilon_{0}a_{0}^{4}}\sqrt{\frac{n^{\prime 2}-1}{n^{3}n^{\prime 5}}}.$ (14)
If a nucleus with Schiff moment $\mathscr{S}_{z}\hat{z}$ interacts with electrons in the state $\left|{\psi}\right\rangle$}, then ${\left\langle{\psi}\right|}V_{cE3}^{\prime}{\left|{\psi}\right\rangle}=-\mathscr{S}_{z}{\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}$. The Schiff moment can be written in terms of the nuclear spin vector $\vec{I}$, as $\mathscr{S}_{z}=\frac{|\vec{\mathscr{S}}|}{|\vec{I}|}I_{z}$, due to the Wigner-Eckart theorem [18]. Thus, when the electron is in state $\left|{\psi}\right\rangle$}, the effective Hamiltonian for the nuclear spin degree of freedom is $H_{\mathrm{NSM}}=-I_{z}\frac{|\vec{\mathscr{S}}|}{|\vec{I}|}{\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}$. When the nuclear spin is in the eigenstate $\left|{I,m_{I}}\right\rangle$}, the Hamiltonian $H_{\mathrm{NSM}}\propto I_{z}$ leads to characteristic energy shifts ${\left\langle{I,m_{I}}\right|}H_{\mathrm{NSM}}{\left|{I,m_{I}}\right\rangle}\propto m_{I}\,{\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}$, which can be measured in experiments (see, e.g., [3, 4]).
Note some interesting properties of $\vec{\mathcal{F}}$. Its expectation value goes to zero if either $c_{s}$ or $c_{p}$ is zero, so only an electron in a superposition of $s$ and $p$ states generates a nonzero $\mathcal{F}$-field in the vicinity of the nucleus. The magnitude of $\vec{\mathcal{F}}$ also scales strongly with the nuclear charge $Z$. These are the reasons why current experiments searching for nuclear Schiff moments use heavy nuclei contained within electrically-polarized atoms or molecules (e.g., [3, 4, 5]), wherein the electrons are in superpositions of atomic $s$ and $p$ states around a nucleus with large $Z$.
In summary, a Schiff moment is a property of a charge distribution that combines aspects of its monopole, dipole, quadrupole and octupole moments. The definition of a nuclear Schiff moment, and the physics of its interaction with electrons in atoms, follow from elementary electrostatics.
Acknowledgments
I have benefited from discussing this topic with Mohit Verma, William Zheng, Wesley Campbell, Andrew Jayich and David DeMille. This work is supported by NSERC and the Canada Research Chairs program.
Appendix: Gradients of $1/r$
Higher-order gradients of the $1/r$ function are rather singular mathematical quantities [15, 16]. To make their behaviour smoother and more physical, let us consider a “regularized” version of $1/r$, which is $\frac{1}{\sqrt{r^{2}+a^{2}}}$. This function is finite and its gradients can be sensibly evaluated, before eventually taking the limit $a\to 0$.
The first-order gradient of $1/r$ can be obtained in a straightforward manner and contains no surprises.
$\lim_{a\to 0}\nabla_{i}\frac{1}{\sqrt{r^{2}+a^{2}}}=-\frac{x_{i}}{(r^{2}+a^{2})^{3/2}}\longrightarrow-\frac{x_{i}}{r^{3}}.$ (15)
The second-order gradient is
$\begin{split}\nabla_{j}\nabla_{i}\frac{1}{\sqrt{r^{2}+a^{2}}}&=-\frac{\delta_{ij}}{(r^{2}+a^{2})^{3/2}}+\frac{3x_{i}x_{j}}{(r^{2}+a^{2})^{5/2}}\\ &=-\frac{\delta_{ij}\,a^{2}}{(r^{2}+a^{2})^{5/2}}+\frac{3x_{i}x_{j}-r^{2}\delta_{ij}}{(r^{2}+a^{2})^{5/2}}.\end{split}$ (16)
As $a\to 0$, the second term tends to $(3x_{i}x_{j}-r^{2}\delta_{ij})/r^{5}$ . But the first term needs some care, especially when $r\ll a$. In this regime, the first term approaches $\delta_{ij}\frac{1}{a^{3}}$. Its volume integral, over any sphere centred at the origin whose radius is $\gg a$, is equal to $\frac{4\pi}{3}{\delta_{ij}}$ regardless of the value of $a$. This fact can be observed from the graph of $\frac{a^{2}}{(r^{2}+a^{2})^{5/2}}$, which is a flat-topped function of width $a$ that falls off rapidly, so that the volume of the sphere over which this function is non-negligible approaches $\frac{4\pi}{3}a^{3}$. Thus its integral over any volume sufficiently larger than $a^{3}$ centred on the origin is equal to $\frac{4\pi}{3}$. In other words,
$\int_{0}^{\infty}\frac{a^{2}}{(r^{2}+a^{2})^{5/2}}\,4\pi r^{2}dr=4\pi\int_{0}^{\infty}\frac{t^{2}\,dt}{(1+t^{2})^{5/2}}=4\pi\left[\frac{t^{3}}{3(1+t^{2})^{3/2}}\right]^{\infty}_{0}=\frac{4\pi}{3}.$ (17)
Since the function $\frac{3}{4\pi}\frac{a^{2}}{(r^{2}+a^{2})^{5/2}}$ only has support over a sphere of radius $\sim a$, and has a volume integral equal to 1 for all values of $a$,
$\lim_{a\to 0}\frac{3}{4\pi a^{3}}\left(1+\frac{r^{2}}{a^{2}}\right)^{-5/2}\to\delta(\vec{r}).$ (18)
Therefore the second-order gradient of $\frac{1}{r}$ is
$\nabla_{j}\nabla_{i}\frac{1}{r}=-\frac{4\pi}{3}\delta_{ij}\,\delta(\vec{r})+\frac{3x_{i}x_{j}-r^{2}\delta_{ij}}{r^{5}}.$ (19)
The first tensor on the right-hand side has nonzero trace, whereas the second tensor is traceless.
The third-order gradient is
$\begin{split}\nabla_{\ell}\nabla_{j}\nabla_{i}\frac{1}{\sqrt{r^{2}+a^{2}}}&=\frac{3\left(\delta_{ij}x_{\ell}+\delta_{j\ell}x_{i}+\delta_{\ell i}x_{j}\right)}{(r^{2}+a^{2})^{5/2}}-\frac{15x_{i}x_{j}x_{\ell}}{(r^{2}+a^{2})^{7/2}}\\ &=\frac{3\left(\delta_{ij}x_{\ell}+\delta_{j\ell}x_{i}+\delta_{\ell i}x_{j}\right)a^{2}}{(r^{2}+a^{2})^{7/2}}+\frac{3\left(\delta_{ij}x_{\ell}+\delta_{j\ell}x_{i}+\delta_{\ell i}x_{j}\right)r^{2}-15x_{i}x_{j}x_{\ell}}{(r^{2}+a^{2})^{7/2}}.\end{split}$ (20)
As $a\to 0$, the second term tends to $\left(3\delta_{\{ij}x_{\ell\}}r^{2}-15x_{i}x_{j}x_{\ell}\right)/r^{7}$ . The first term again needs some care. From Equation (18),
$\lim_{a\to 0}-\frac{15}{4\pi a^{5}}\,x_{\ell}\,\left(1+\frac{r^{2}}{a^{2}}\right)^{-7/2}\to\nabla_{\ell}\delta(\vec{r}),$ (21)
so that the first term in Equation (20) tends to $\frac{4\pi}{5}\delta_{\{ij}\nabla_{\ell\}}\delta(\vec{r})$. Therefore
$\nabla_{\ell}\nabla_{j}\nabla_{i}\frac{1}{r}=-\frac{4\pi}{5}\delta_{\{ij}\nabla_{\ell\}}\delta(\vec{r})+\frac{3\delta_{\{ij}x_{\ell\}}r^{2}-15x_{i}x_{j}x_{\ell}}{r^{7}}.$ (22)
Once again, the first tensor on the right-hand side has nonzero trace, whereas the second tensor is traceless.